40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 11 01:05:29 2011, kiwamu, Update, LSC, obtained Xarm PDH signal 
    Reply  Wed Jan 12 01:39:49 2011, kiwamu, Update, LSC, locked X arm 
Message ID: 4141     Entry time: Wed Jan 12 01:39:49 2011     In reply to: 4131
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: locked X arm 

[Suresh, Kiwamu]

 We eventually succeeded in locking X arm with the infrared beam.

The PDH signal is taken at MCL's ADC instead of c1lsc's, and fedback to MC2_POS through the MCL path.

Right now the lock is not so stable for some reasons, so we need to investigate it more.

 

(what we did)

 - strung a long BNC cable to connect the demodulated signal and the ADC of c1ioo.

          We didn't touch anything on the demodulation system, so the setup for the demodulation is exactly the same as that of yesterday (see here).

 - disconnected the actual MCL cable from the ADC breakout board at 1X2 rack. And put the demodulated signal onto it.

 - checked the analog dewhitening filter state for the MC2 coil driver, found the analog filter are always off.

  So we just made simDW and invDW always on.

- changed the gain of the MCL loop to have a stable lock for the X arm.

    right now a reasonable setup in the MCL filters are:

         FM1:ON, FM10:ON, G=0.1

 - In fact the lock of the MC is not so stable compared to before, frequently an attempt of locking the X arm leads to the unlock of the MC.

ELOG V3.1.3-